OpenStudy (anonymous):

A snare drum has diameter fourteen inches. A bass drum has diameter thirty-six inches. What is the difference in the areas of the two drums? sixty-nine point zero eight square inches four hundred eighty-four square inches eight hundred sixty-three point five square inches three thousand nine hundred fifteen point five eight square inches I will medal and fan

OpenStudy (welshfella):

area = pi r^2 where r = radius radius = 1/2 the diameter

OpenStudy (anonymous):

so wat is the answer?

OpenStudy (mathstudent55):

That is for you to calculate.

OpenStudy (mathstudent55):

Do you know how to find the area of a circle?

OpenStudy (anonymous):

nope

OpenStudy (retireed):

AreaBassDrum - AreaSnareDrum = A = pi r^2 = pi*[d(Bass)/2]^2 - pi*[d(snare)/2]^2

OpenStudy (mathstudent55):

You need the formula above and the radius. The radius is half of the diameter.

OpenStudy (mathstudent55):

If a circle has diameter 14 inches, what is its radius?

OpenStudy (anonymous):

do u guys know the answer but u r just telling me step by step?

OpenStudy (anonymous):

7 r

OpenStudy (mathstudent55):

I don't know the answer because I have not solved this problem. I am trying to guide you in the solution of the problem by showing you each step you need to do.

OpenStudy (anonymous):

k thx i just had to knw

OpenStudy (anonymous):

so is it 7r?

OpenStudy (mathstudent55):

Good. One radius is 7 in. The other circle has diameter 36 inches. What is its radius?

OpenStudy (anonymous):

oh ok good

OpenStudy (anonymous):

18 r

OpenStudy (mathstudent55):

Good. We have two circles with radii 7 inches and 18 inches. Now we find the two areas. The formula for the area of a circle is \(A = \pi r^2\) Area equals "pi" times the radius squared. The radius squared is the radius times the radius. pi is a Greek letter used in math that has a value of approximately 3.14

OpenStudy (mathstudent55):